1.   Why do you dilute the fibroblasts in trypan blue before counting them?  What will the trypan blue tell you about the cells?

2.   Why do you add trypsin to the fibroblasts in the culture dish? How does trypsin do this?

3.   You have counted the 4 corner squares of a hemocytometer grid (each square is 1mm2 and the distance of the coverslip above the squares is 0.1mm) and have come up with a total cell count of 243 cells. What is your original cell concentration? Use appropriate units. The dilution factor you should use is 2.

4.   If in the previous question, 212 of the cells were NOT stained with trypan blue, what would be the percent viability of the cells?

5.   If you added 250µl of fibroblasts at a concentration of 1.75 x 105 cells/ml to a new culture dish that already contained 2.250 ml of tissue culture media, what would the concentration be of the cells in the new dish?  

6.   In the FluorMicros lab, 12-well plates, each containing 1ml of cell culture media, received control solutions or ceramide. In well C4, 3µl of a 2mM ceramide solution was added to the cells. What was the final concentration of ceramide for that particular treatment?

                        A.         2µM

                        B.         6µM

                        C.         2mM

                        D.         6mM

                        E.         This cannot be determined from the given information.
Answer #1

1. Tryphan blue is a dye used to test the viability of a cell. Live cells have an integrated cell membrane, so they dont allow any dye to pass through the membrane. Therefore, the live cells will have a clear cytoplasm and a dead one will take up the colour. By diluting the cells with tryphan blue we can count the viable cells.

2. Trypsin is a protein degrading enzyme which cuts the polypeptide after lysine. Trypsin breaks the protein which helps the cells to adhere on the culture dish. Therefore, it dissociates the cells from the culture plate.

3. Cell count in hemeocytometer is calculated by the following formula

Number of cells = {(number of cells counted) (dilution factor)/ (number of squares counted) (volume of one square)} X 10000

Volume of liquid occupies in each square of 1mm2 X 0.1mm is 100nl.

Number of squares counted = 4

Number of counted cells = 243

Dilution factor = 2

= {(243 X 2)/ (4 X 100nl)} X 10000 = (486/ 400nl) X 10000 = 1.215nl X 10000 = 12150 cells

4. If 212 cells were not stained then, 212 cells are alive.

Number of viable cells = {(212X2)/ (4X100nl)} X 10000 = (424/ 400nl) X 10000 = 1.06 X 10000

= 10600 cells

Therefore percentage of viable cells = (10600/12150) X 100 = 87.24%